In an adiabatic change, the pressure and temperature of a monoatomic gas are related to Pdirectly proposional to T^c,
where c equals

Grade:11

1 Answers

Arun
5 years ago
Dear Sarthak
 
As we know
 
PV^x = constant
 
Now we have to find relation between P and T, hence V = nRT/P
 
Here n and R constant. hence V = T/P
 
P (T/P)^x = constant
 
P^(1-x) T^x = constant
 
Hence
 
P = T( – x/1-x)
 
hence c = – x/(1 – x)